Payton Company has the following segment revenues for the two most recent fiscal years.

Current Year (in millions) Prior Year (in millions)
China $775.00 $650.00
Canada 325.50 245.50
Other countries 215.50 168.50
Total revenues $1,316.00 $1,064.00

Required:

Prepare a vertical analysis of the segment data. Round your percentage value to one decimal place.

Prepare a vertical analysis of the segment data. Round your percentage value to one decimal place.

1

Current Year (in millions)

Percent

Prior Year (in millions)

Percent

2

China

$775.00

$650.00

3

Canada

325.50

245.50

4

Other countries

215.50

168.50

5

Total revenues

$1,316.00

$1,064.00
Answer #1
Current year Prior year
China (775/1316) = 58.9% (650/1064) = 61.1%
Canada (325.5/1316) = 24.7% (245.5/1064) = 23.1%
Other (215.5/1316) = 16.4% (168.5/1064) = 15.8%
Total 100% 100%

Comment if you face any issues